OpenStudy (anonymous):

Please help! Solve the following system using any method: x+y=2 x^2-y=0 Another solution is: a. (1,1) b. (-1,1) c. (-1,-1) d. (1, -1)

OpenStudy (anonymous):

from 2nd equation y=x^2 put this value of y in 1st equation x+x^2=2 x^2+x=2 x^2+x-2=0 x^2+2x-x-2=0 x(x+2)-1(x+2)=0 (x-1)(x+2)=0 x=1,x=-2

OpenStudy (anonymous):

when x=1,y=1 and when x=-2,y=4

OpenStudy (anonymous):

so A is the answer

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

From first equation: x = 2 - y Put this in second equation: (2 - y)^2 - y = 0 \[y^2 - 4y + 4 -y = 0\] \[y^2 - 5y +4 = 0\] \[(y-4)(y-1) = 0\] y = 4 or y = 1 So, from this x = 1 when y = 1 and x = -2 when y = 4 So, the required answer is : (1, 1) i.e, A..

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Welcome dear...